# Challenges for first-principles nuclear structure: 11Li and 29F

### v1

@@ Abstract
Ab initio calculations of atomic nuclei have had many successes in recent years. Nonetheless, important challenges that resist even brute-force calculation remain. As archetypal examples of these challenges, we consider 11Li and 29F, well known halo nuclides situated on islands of inversion. The deformed intruder levels, which are primarily two-particle, two-hole neutron excitations with respect to naive spherical shell model configurations, are slow, with respect to increases of the model space, to take their rightful place as the lowest levels; in fact the spherical and deformed configurations likely mix strongly. We suggest these systems prototype the challenges for other important intruder states, and can serve as useful testbeds for potential approaches.
@@ Introduction.
Simple models can be challenged by messy realities; adapting to reality can lead to new, more insightful models. As an early example of this, the spherical independent-particle shell-model of atomic nuclei was confronted by the reality of collective quadrupolar dynamics, which lead, first to the deformed Nilsson model [1, 2], and then to Elliott’s SU(3) picture [3, 4, 5, 6, 7, 8]. Particularly instructive is the history of the configuration-interaction (CI) shell-model, where one expands a many-body state in a basis, allowing the Schrödinger equation to be written as a matrix eigenvalue problem [9].
@@
Because of the strength of the nuclear force, one must take account of configurations that produce holes in the core, and those that catapult particles into the excluded orbitals. Originally this was done through perturbation theory [9, 12]. But careful analyses showed that such perturbative calculations would generically run into difficulties [13, 14].
To explain the difficulties, one must introduce the enveloping concept of intruders. In the simplest definition of intruders, one divides the infinite Hilbert space into the model space and the excluded space, represented by projection operators $P$ and $Q$, respectively, such that $P+Q=1$. Ideally, low-lying wave functions live mostly in the model space $P$; wave functions in the excluded space $Q$ are expected to lie higher in energy. An intruder upends this expectation: it is a wave function lying low in the spectrum, but dominated by components from the excluded space.
The existence of intruder states in nuclei was recognized early on; for example the realization [15] that the ground state of 11Be has $J^{\pi}=1/2^{+}$, whereas the naive shell model would suggest instead $1/2^{-}$. Some of best known intruders may be the primarily four-particle, four-hole (alpha cluster excitation) Hoyle state in 12[16, 17] and the related first excited $0^{+}$ state at 6.05 MeV in 16[18, 19], and $0^{+}$ states at 3.35 and 5.21 MeV in 40Ca [20]. The notorious islands of inversion [21, 22], where the ground state is itself an intruder relative to the valence space, as in 11Be  [15], provide yet more examples. Intruder states are often associated with shape coexistence, where spherical and deformed states lie close in energy and mix. The coexisting deformed states can be interpreted in terms of particle-hole excitations across shell boundaries and are found in many regions of the chart of the nuclides; see [23], especially their Figs. 2 and 8.
The term “intruder” implies something unnatural or unusual about a state, especially when contrasted against so-called “normal” states as we do below. Yet it can be argued that while these deformed particle-hole states are intruders from the viewpoint of spherical orbitals, they are in fact are naturally expected by applying the aufbau principle to a basis of deformed single-particle orbitals such as the Nilsson picture [24, 25, 26, 27]. Indeed, soon after it was suggested that 11Li has a neutron halo, it was pointed out this could be easily explained in the Nilsson model [24].
@@
It was the introduction of the no-core shell model [9, 31] (NCSM) that revived nuclear configuration-interaction as a potent tool with a rigorous ab initio basis. Because the NCSM from the very beginning includes, in principle, all excitations, and those excitations are handled nonperturbatively, the original intruder crisis went away (or, to put it another way, since there is no core, there can be no particle-hole excitations to cause trouble).
No-core shell model basis states themselves are constructed from harmonic oscillator single-particle states described by the oscillator frequency $\omega$, or more commonly, the oscillator energy parameter $\hbar\omega$. A key label of many-body configurations is the number $N_{\text{ex}}$} of harmonic oscillator excitation quanta relative to the lowest filling of oscillator shells. Configurations of a given $N_{\text{ex}}$} are sometimes termed $N_{\mathrm{ex}}\hbar\omega$ configurations: the lowest configurations, with $N_{\mathrm{ex}}=0$, are $0\hbar\omega$ configurations, etc. A given model space is then labeled by $N_{\mathrm{max}}$, the maximum excitation allowed, or, alternately, as an $N_{\mathrm{max}}\hbar\omega$ space. Thus, while there is formally no valence space in the NCSM, it is nonetheless convenient to refer to the $N_{\mathrm{max}}=0$ or $0\hbar\omega$ space as the naive valence space. Then we define states dominated by configurations in the naive valence space as “normal” states, and states dominated by configurations outside the naive valence space, in this scheme, as “intruders.” Applying the concept of intruders to the NCSM is not new [32, 33, 34, 35, *caprio2022:10be-shape-sdanca21, 37, 38, 39].
Figure 1: Schematic of normal (left side) versus two-particle, two-hole intruder (right side) configurations. Dashed lines represent spherical shell closures. For our example of 11Li, the filled proton orbitals (diagonal fill) is $0s$ while the filled neutron orbitals (solid fill) consist of 0s0s-$0p$. For 29F, the filled proton and neutron orbitals are 0s0s-$0p$ and 0s0s-0p0p-$1s0d$, respectively.
In this paper, we discuss detailed NCSM calculations of two cases from the islands of inversion, 11Li and 29[22, 40]. These two exemplars are parallel, in that in the naive spherical shell model, one would expect the lowest configuration to consist of filled neutron oscillator shells ($0s$ and $0p$ in the case of 11Li, and $0s$, $0p$, and $1s0d$ for 29F) and a single proton outside a filled shell. Experimental evidence, however, supports the ground state having a significant contribution from two-particle, two-hole neutron excitations. Fig. 1 provides a schematic illustration. Our calculations highlight significant differences between the normal and intruder states: the normal states are nearly spherical, while the intruder states are strongly deformed.
@@
We use two sets of interaction matrix elements. The first is derived from the Entem-Machleidt next-to-next-to-next-to-leading-order (N3LO) chiral effective theory [41], softened by similarity-group renormalization [42, 43] (SRG) evolution to a resolution parameter value of $\lambda=2.0\,{\mathrm{fm}}^{-1}$. We do not include three-body forces. The second interaction, Daejeon16 [44], is phase-shift-equivalent to the Entem-Machleidt interaction, but adjusted to reproduce the binding energies (and some excitation energies) in several nuclei with $A\leq 16$, without a three-body force. The latter is known to be “softer” than the former, since it has been more drastically SRG evolved, to $\lambda=1.5\,{\mathrm{fm}}^{-1}$. This leads to faster convergence of energies with increase of the model space, as we see in our calculations. For our observables such as the quadrupole moment and radius, we use the bare (unrenormalized) operators and charges. We choose the value of $\hbar\omega$ for each nuclide and interaction by approximately minimizing the lowest energy (ostensibly the ground state energy, although in our calculations the lowest state is predominantly “normal”, while the true ground state would in fact be predominantly “intruder”). NCSM calculations were carried out using the codes BIGSTICK [45] and MFDn [46, *shao2018:ncci-preconditioned].
To gain insight into the structure of the wave functions, we decompose via the Lanczos algorithm [48, 49, 50, 36], into irreps of Elliott’s SU(3) [51, 36, 39]. These irreps can be in turn approximately interpreted in terms of deformation parameters $\beta$ and $\gamma$ [52]. We can also simultaneously decompose into $N_{\mathrm{ex}}$, the number of oscillator excitations.
@@
Figure 2: (Left) Relative energies for the first two $1/2^{-}$ and $3/2^{-}$ levels, and first $5/2^{-}$ and $7/2^{-}$ levels, of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}, as a function of $N_{\mathrm{max}}$, for (a) a chiral N3LO interaction and (b) the Daejeon16 interaction. Calculated values are shown as functions of $N_{\text{max}}$}, at fixed basis parameter of ${\hbar\omega}=18\,{\mathrm{MeV}}$ for N3LO and ${\hbar\omega}=12.5\,{\mathrm{MeV}}$ for Daejeon16. The shading of the symbols indicates the structure, normal (shaded) or intruder (open), which might naively be expected for the level given the energy evolution, although the first two $1/2^{-}$ levels are undergoing an avoided crossing at higher $N_{\text{max}}$} in the Daejeon16 calculations (see text). (Right) Decompositions by $N_{\text{ex}}$} for first two $3/2^{-}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}, in a calculation with ${N_{\text{max}}}=8$, for the Daejeon16 interaction (with ${\hbar\omega}=12.5\,{\mathrm{MeV}}$).
@@
11Li was an early and dramatic exemplar of halo nuclei [53]. Neutron knockout experiments show nearly equal contributions of neutron $1s_{1/2}$ and $0p_{1/2}$ orbitals [54], providing strong evidence for a ground state comprised of a nearly equal mix of normal and intruder states. 11Li is well known to be a challenge to describe fully microscopically, i.e., in terms of three protons and eight neutrons, for example in the NCSM [55, 56, 57], or using fermionic molecular dynamics (FMD) [58]. The FMD calculation addresses mixing of $(p_{1/2})^{2}$ and $(s_{1/2})^{2}$ configurations, that is, normal and intruder states. In many other calculations, 11Li has been approximated as a three-body system, typically $^{9}\mathrm{Li}+2n$ [59, 60, 61, 62, 63, 64, 65, 66, 67, 68], also allowing $(p_{1/2})^{2}$-$(s_{1/2})^{2}$ mixing to be investigated.
@@
Our calculated excitation spectrum is shown in Fig. 2, as a function of $N_{\mathrm{max}}$ for both the chiral N3LO and Daejeon16 interactions (at fixed basis parameter ${\hbar\omega}=18\,{\mathrm{MeV}}$ and $12.5\,{\mathrm{MeV}}$, respectively). In the rightmost panels, we display the $N_{\text{ex}}$} decomposition of the ground state and excited $3/2^{-}$ states, in particular, illustrating how we may identify normal and intruder states from their NCSM wave functions [35, 37, 39].
@@
The $0{\hbar\omega}$ shell model picture for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} gives just $3/2^{-}$ and $1/2^{-}$ states, corresponding to finding the valence neutron in the $p_{3/2}$ and $p_{1/2}$ orbitals, respectively. Indeed, the two lowest calculated states (shaded symbols in Fig. 2) match this expectation, and, from their decomposition into configurations with different $N_{\text{ex}}$}, they are seen to be normal (with an exception related to mixing at high $N_{\text{max}}$}, as discussed below). This is illustrated for the $3/2^{-}$ ground state in Fig. 2 (bottom right). Namely, the single largest contribution to the probability comes from ${N_{\text{ex}}}=0$ (in this case, a single configuration, the filled neutron $0p$ shell), although, naturally for a no-core calculation (see, e.g., Fig. 8 of Ref. [37]) this $0{\hbar\omega}$ configuration is “dressed” with a tail of smaller contributions from excited configurations.
@@
Then, several “intruder” states arise at higher energy in the calculation (open symbols in Fig. 2), starting with a closely spaced $3/2^{-}$ and $1/2^{-}$. (There are many more such states, of which we only show the first $7/2^{-}$ and $5/2^{-}$ as representative examples.) While these start at high energy, for low $N_{\text{max}}$}, they rapidly descend with increasing $N_{\text{max}}$}. The intruder nature of the excited $3/2^{-}$ and $1/2^{-}$ states (again, with an exception related to mixing, as discussed below) is illustrated for the excited $3/2^{-}$ state in Fig. 2 (top right). Note the negligible $0{\hbar\omega}$ contribution, of a few percent, while the strongest contribution comes from $2{\hbar\omega}$ configurations, which are again dressed with a tail of more highly excited oscillator configurations.
@@
Because Daejeon16 is a “softer” interaction, it is perhaps unsurprising that the intruder states found in the Daejeon16 calculation [Fig. 2(b)] descend in energy much sooner (with increasing $N_{\mathrm{max}}$) than those found in the N3LO calculation [Fig. 2(a)]. They also start to level off in energy, suggesting they are approaching the values they would converge to in an untruncated calculation.
@@
As the energies of the intruder states approach those of the normal states, of the same angular momentum and parity, these states can mix. In the Daejeon16 calculations [Fig. 2(b)], the normal and intruder $1/2^{-}$ states (upward triangles) approach in energy first (due to the high excitation energy of the normal $1/2$- state, relative to the spacing between the intruder $3/2^{-}$ and $1/2^{-}$ states). They approach to within $\sim 2\,{\mathrm{MeV}}$, with increasing $N_{\text{max}}$}, but then get no closer. The excited $1/2^{-}$ state (open upward triangle) initially tracks the other intruder states, at lower $N_{\text{max}}$}, but then appears to hit a floor, at lower $N_{\text{max}}$}, while the other intruder states continue to fall in energy. On the other hand, the first $1/2^{-}$ state (shaded upward triangle), while stable in excitation energy below ${N_{\text{max}}}=4$, is pushed down in excitation energy, for higher $N_{\text{max}}$}. This behavior is suggestive of the level repulsion arising in an avoided crossing (e.g., Ref. [69]). Indeed, the $N_{\text{ex}}$} decompositions (not shown) of the wave functions for the calculated $1/2^{-}$ states suggest that the “normal” and “intruder” states are nearly maximally mixed by ${N_{\text{max}}}=12$. (The mixing matrix element between the normal and intruder states can thus be deduced to be on the order of half the minimal separation, or $\sim 1\,{\mathrm{MeV}}$.) With further increasing $N_{\text{max}}$}, we may expect that the intruder $1/2^{-}$ would emerge from the other side of the avoided crossing and resume tracking the energies of the other intruder states, thus supplanting the normal $1/2^{-}$ state as the lowest $1/2^{-}$ state.
Although the normal and intruder $3/2^{-}$ states [squares in Fig. 2(b)] do not yet approach as closely at the $1/2^{-}$ states in these calculations, the $N_{\text{ex}}$} decompositions of the wave functions (not shown) suggest about $10\%$ mixing already by ${N_{\text{max}}}=12$. We can anticipate that the mixing will continue to increase, as the intruder state continues to approach the ground state. How strongly the states will mix, and whether or not they might actually cross, depends in detail on the energy convergence and the strength of the mixing matrix element for the $3/2^{-}$ states, so we do not attempt a quantitative analysis here. We simply wish to highlight the importance of the intruder state and mixing in understanding even the ground state of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} in an ab initio calculation.

### v2

@@ Abstract
Ab initio calculations of atomic nuclei have had many successes in recent years. Nonetheless, important challenges that resist even brute-force calculation remain. As archetypal examples of these challenges, we consider 11Li and 29F, well known halo nuclides situated on islands of inversion. The deformed intruder levels, which are primarily two-particle, two-hole neutron excitations with respect to naive spherical shell model configurations, are slow, with respect to increases of the model space, to take their rightful place among, and potentially mix with, the lowest levels. We suggest these systems prototype the challenges for other important intruder states, and can serve as useful testbeds for potential approaches.
@@ Introduction.
Simple models can be challenged by messy realities; adapting to reality can lead to new, more insightful models. As an early example of this, the spherical independent-particle shell-model of nuclei was confronted by collective quadrupolar dynamics, which lead, first to the deformed Nilsson model [1, 2], and then to Elliott’s $\mathrm{SU}(3)$ picture [3, 4, 5, 6, 7, 8]. Particularly instructive is the history of the configuration-interaction shell-model, where one expands a many-body state in a basis, allowing the Schrödinger equation to be written as a matrix eigenvalue problem [9].
@@
Because of the strength of the nuclear force, one must take account of those configurations with holes in the core, and those with particles catapulted into the excluded orbitals. Originally this was done through perturbation theory [9, 12]. But careful analyses showed that such perturbative calculations would generically run into difficulties [13, 14].
To explain the difficulties, one must introduce the concept of intruders. In the simplest definition of intruders, one divides the infinite Hilbert space into the model space and the excluded space, represented by projection operators $P$ and $Q$, respectively, such that $P+Q=1$. Ideally, low-lying wave functions live mostly in the model space $P$; wave functions in the excluded space $Q$ are expected to lie higher in energy. An intruder upends this expectation: a state lying low in the spectrum, but dominated by components from the excluded space.
The existence of intruder states in nuclei was recognized early on: for example the realization [15] that the ground state of 11Be has $J^{\pi}=1/2^{+}$, whereas the naive shell model would suggest $1/2^{-}$. Some of the best known intruders are the four-particle, four-hole (alpha cluster excitation) Hoyle state in 12[16, 17] and the related first excited $0^{+}$ state at 6.05 MeV in 16[18, 19], and $0^{+}$ states at 3.35 and 5.21 MeV in 40Ca [20]. The notorious islands of inversion [21, 22], where the ground state is itself an intruder relative to the valence space, as in 11Be [15], provide more examples. Intruder states are often associated with shape coexistence, where spherical and deformed states lie close in energy. The coexisting deformed states can be interpreted in terms of particle-hole excitations across shell boundaries and are found in many regions of the chart of the nuclides (see Ref. [23], especially their Figs. 2 and 8).
The term “intruder” implies something unnatural or unusual about a state, especially when contrasted against so-called “normal” states as we do below. Yet while these deformed particle-hole states are intruders from the viewpoint of spherical orbitals, they are in fact are naturally expected by applying the aufbau principle to a basis of deformed single-particle orbitals such as the Nilsson picture [24, 25, 26, 27]. Indeed, soon after it was suggested that 11Li has a neutron halo, it was pointed out this could be easily explained in the Nilsson model [24].
@@
It was the introduction of the no-core shell model [9, 31] (NCSM) that revived nuclear configuration-interaction as a potent tool with a rigorous ab initio basis. Because the NCSM from the very beginning includes, in principle, all excitations, and those excitations are handled nonperturbatively, the original intruder crisis went away (or rather, since there is no core, there can be no particle-hole excitations to cause trouble).
No-core shell model basis states themselves are constructed from harmonic oscillator single-particle states described by the oscillator frequency $\omega$, or more commonly, the oscillator energy parameter $\hbar\omega$. A key label of many-body configurations is the number $N_{\text{ex}}$} of harmonic oscillator excitation quanta relative to the lowest filling of oscillator shells. Configurations of a given $N_{\text{ex}}$} are sometimes termed $N_{\mathrm{ex}}\hbar\omega$ configurations: the lowest configurations, with $N_{\mathrm{ex}}=0$, are $0\hbar\omega$ configurations, etc. A given model space is then labeled by $N_{\mathrm{max}}$, the maximum excitation allowed, or, alternately, as an $N_{\mathrm{max}}\hbar\omega$ space. Thus, while there is formally no valence space in the NCSM, it is nonetheless convenient to refer to the $N_{\mathrm{max}}=0$ or $0\hbar\omega$ space as the naive valence space. Then we define states dominated by configurations in the naive valence space as “normal” states, and states dominated by configurations outside the naive valence space, in this scheme, as “intruders.” Applying the concept of intruders to the NCSM is not new [32, 33, 34, 35, 36, 37, 38, 39].
Figure 1: Schematic of normal (left side) versus two-particle, two-hole intruder (right side) configurations. Dashed lines represent spherical shell closures. For our example of 11Li, the filled proton orbital (hashed fill) is $0s$ while the filled neutron orbitals (solid fill) consist of 0s0s-$0p$. For 29F, the filled proton and neutron orbitals are 0s0s-$0p$ and 0s0s-0p0p-$1s0d$, respectively.
In this paper, we discuss detailed NCSM calculations of two cases from the islands of inversion, 11Li and 29[22, 40, 41]. These two exemplars are parallel, in that in the naive spherical shell model, one would expect the lowest configuration to consist of filled neutron oscillator shells ($0s$ and $0p$ in the case of 11Li, and $0s$, $0p$, and $1s0d$ for 29F) and a single proton outside a filled shell. Experimental evidence, however, supports the ground state having a significant contribution from two-particle, two-hole neutron excitations. Fig. 1 provides a schematic illustration. Our calculations highlight significant differences between the normal and intruder states: the normal states are nearly spherical, while the intruder states are strongly deformed.
@@
We use two sets of interaction matrix elements. The first is derived from the Entem-Machleidt next-to-next-to-next-to-leading-order (N3LO) chiral effective theory [42], softened by similarity-group renormalization [43, 44] (SRG) evolution to a resolution parameter value of $\lambda=2.0\,{\mathrm{fm}}^{-1}$. We do not include three-body forces. The second interaction, Daejeon16 [45], is again derived from the Entem-Machleidt interaction, but in this case more drastically softened by SRG evolution, to $\lambda=1.5\,{\mathrm{fm}}^{-1}$ (after which phase-shift-equivalent transformations are applied to reproduce the binding energies, and selected excitation energies, in several nuclei with $A\leq 16$, without including any induced three-body force). The resulting “softer” interaction leads to faster convergence of energies with increase of the model space, as we see in our calculations. For our observables, namely, the quadrupole moment and radius, we use the bare (unrenormalized) operators and charges.
We choose the value of $\hbar\omega$ for each nuclide and interaction by approximately minimizing the lowest energy (ostensibly the ground state energy, although in our calculations the lowest state is predominantly normal, while the true ground state would in fact be predominantly intruder). NCSM calculations were carried out using the codes BIGSTICK [46] and MFDn [47, 48].
To gain insight into the structure of the wave functions, we decompose them, via the Lanczos algorithm [49, 50, 51, 36], into irreps of Elliott’s $\mathrm{SU}(3)$ [51, 36, 39]. These irreps can be approximately interpreted in terms of deformation parameters $\beta$ and $\gamma$ [52]. We also simultaneously decompose into $N_{\mathrm{ex}}$, the number of oscillator excitations.
Figure 2: Relative energies for the normal $1/2^{-}$ and $3/2^{-}$ (blue, shaded symbols) and first intruder $1/2^{-}$ through $7/2^{-}$ (red, open symbols) levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}, for (a) a chiral N3LO interaction and (b) the Daejeon16 interaction, shown as functions of $N_{\text{max}}$} (at fixed $\hbar\omega$}, as indicated). Although states are designated as normal or intruder in this figure according to what might naively be expected for the level given the energy evolution, the first two $1/2^{-}$ levels start to undergo an avoided crossing at higher $N_{\text{max}}$} in the Daejeon16 calculations (see text).
@@
Figure 3: Decompositions by $N_{\text{ex}}$} (left) and by $\mathrm{SU}(3)$ jointly with $N_{\text{ex}}$} (right), for the normal (bottom) and intruder (top) $3/2^{-}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}, for the Daejeon16 interaction, in calculations with $N_{\mathrm{max}}=8$ and ${\hbar\omega}=12.5\,{\mathrm{MeV}}$. The $\mathrm{SU}(3)$ decompositions are shown arranged by the Bohr deformation variables corresponding to the given $\mathrm{SU}(3)$ quantum numbers, and include contributions from the $0{\hbar\omega}$ (blue, shaded circles) and $2{\hbar\omega}$ (red, open circles) spaces. Contributions from irreps which are degenerate with respect to the $\mathrm{SU}(3)$ Casimir operator (connected by dotted lines) cannot be distinguished (and, for plotting purposes, such contributions have, arbitrarily, been distributed equally between these irreps).
@@
The nucleus 11Li provided an early and dramatic exemplar of halo nuclei [53]. Neutron knockout experiments show nearly equal contributions of neutron $1s_{1/2}$ and $0p_{1/2}$ orbitals [54], providing strong evidence for a ground state comprised of a nearly equal mix of normal and intruder states. 11Li is well known to be a challenge to describe fully microscopically, i.e., in terms of three protons and eight neutrons, for example in the NCSM [55, 56, 57], or using fermionic molecular dynamics (FMD) [58]. The FMD calculation addresses mixing of $(p_{1/2})^{2}$ and $(s_{1/2})^{2}$ configurations, that is, normal and intruder states. The nucleus 11Li has also been approximated as a three-body system, typically $^{9}\mathrm{Li}+2n$ [59, 60, 61, 62, 63, 64, 65, 66, 67, 68], allowing $(p_{1/2})^{2}$-$(s_{1/2})^{2}$ mixing to be investigated.
@@
Our calculated excitation spectrum is shown in Fig. 2, as a function of $N_{\mathrm{max}}$, for both the chiral N3LO and Daejeon16 interactions (at fixed basis parameter ${\hbar\omega}=18\,{\mathrm{MeV}}$ and $12.5\,{\mathrm{MeV}}$, respectively). This figure highlights the different evolution with $N_{\text{max}}$} for normal (blue, shaded symbols) and intruder (red, open symbols) states. An illustration of how we may identify normal and intruder states from their NCSM wave functions [35, 37, 39] is provided in Fig. 3 (left), where we display the decomposition of the wave functions into ${N_{\text{ex}}}{\hbar\omega}$ components, in particular, for the $3/2^{-}$ levels.
The $0{\hbar\omega}$ shell model picture for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} gives just $3/2^{-}$ and $1/2^{-}$ states, corresponding to finding the valence proton in the $p_{3/2}$ and $p_{1/2}$ orbitals, respectively. Indeed, the two lowest calculated states (blue, shaded symbols in Fig. 2) match this expectation.
@@
From their decomposition into configurations with different $N_{\text{ex}}$}, these states are clearly normal, as illustrated for the $3/2^{-}$ ground state in Fig. 3 (bottom left). Namely, the single largest contribution to the probability comes from ${N_{\text{ex}}}=0$ (in this case, a single configuration, the filled neutron $0p$ shell), although, naturally for a no-core calculation (see, e.g., Fig. 8 of Ref. [37]), this $0{\hbar\omega}$ configuration is “dressed” with a tail of smaller contributions from excited configurations.
@@
Then, several intruder states arise at higher energy in the calculation (red, open symbols in Fig. 2), starting with a closely spaced $3/2^{-}$ and $1/2^{-}$. (There are many more such states, of which we only show the first $7/2^{-}$ and $5/2^{-}$ as representative examples.) While these start at high energy, for low $N_{\text{max}}$}, they rapidly descend with increasing $N_{\text{max}}$}.
The intruder nature of the excited $3/2^{-}$ and $1/2^{-}$ states is illustrated for the excited $3/2^{-}$ state in Fig. 2 (top left). Note the negligible $0{\hbar\omega}$ contribution, of a few percent, while the strongest contribution comes from $2{\hbar\omega}$ configurations, which are again dressed with a tail of more highly excited oscillator configurations.
@@
Because Daejeon16 is a “softer” interaction, it is perhaps unsurprising that the intruder states found in the Daejeon16 calculation [Fig. 2(b)] descend in energy much sooner (with increasing $N_{\mathrm{max}}$) than those found in the N3LO calculation [Fig. 2(a)]. They also start to level off in energy, suggesting they are approaching the values they would attain in an untruncated calculation.
@@
As an intruder state approaches the normal state of the same angular momentum and parity, these states can mix. In the Daejeon16 calculations [Fig. 2(b)], it is the $1/2^{-}$ normal and intruder states (upward triangles) which approach in energy first, due to the high excitation energy of the normal $1/2^{-}$ state (and relatively smaller spacing between the intruder $3/2^{-}$ and $1/2^{-}$ states). The excited $1/2^{-}$ state initially tracks the other intruder states in energy, but then appears to hit a floor at higher $N_{\text{max}}$}, while the other intruder states continue to fall in energy. On the other hand, the normal $1/2^{-}$ state is pushed down in excitation energy for higher $N_{\text{max}}$}. The two levels approach to within $\approx 2\,{\mathrm{MeV}}$, but no closer. This behavior is suggestive of the level repulsion arising in an avoided crossing [69, 70]. Such a mixing picture is borne out, e.g., by examination of $N_{\text{ex}}$} decompositions (similar to those in Fig. 3, but not shown), which suggest that the normal and intruder states $1/2^{-}$ are nearly maximally mixed by ${N_{\text{max}}}=12$. Although the normal and intruder $3/2^{-}$ states (squares) do not yet approach as closely at the $1/2^{-}$ states in these calculations, we can anticipate that they will likewise strongly mix, as the intruder state $3/2^{-}$ continues to approach the normal $3/2^{-}$ “ground state”.